# Enhancing LLM Evaluations: The Garbling Trick

### v2

@@ Abstract
As large language models (LLMs) become increasingly powerful, traditional evaluation metrics tend to saturate, making it challenging to distinguish between models based on their performance. We propose a general method to transform existing LLM evaluations into a series of progressively more difficult tasks. These enhanced evaluations emphasize reasoning capabilities and can reveal relative performance differences that are not apparent in the original assessments.
To demonstrate the effectiveness of our approach, we create a new multiple-choice test corpus, extend it into a family of evaluations, and assess a collection of LLMs. Our results offer insights into the comparative reasoning abilities of these models, particularly highlighting distinctions between OpenAI’s o1-preview and Google’s gemini-pro-1.5-002.
@@ 1 Introduction
In this paper, we present a straightforward method for transforming any text-based evaluation into a family of increasingly challenging tests. This is a meta-evaluation technique: it takes an existing evaluation framework and converts it into a new one. To demonstrate our method, we generate a new evaluation framework and then extend it into a harder family of tests. Even this simple example reveals non-obvious differences in reasoning performance among models.
@@ 2 The Garbling Trick
Evaluations of language models have shifted from testing syntax (e.g., model perplexity) to assessing semantics (e.g., answering multiple-choice questions) [8]. The garbling trick enables us to interpolate between these extremes and examine both aspects simultaneously. As the garbling rate increases, the LLM faces two additional tasks not present in the original evaluation. First, the LLM must infer as much as possible about the underlying text from its garbled version. Second, given that it can only partially or uncertainly recover the text, the LLM must answer the question with incomplete information. Selecting the correct answer thus requires the LLM to engage in multiple forms of reasoning concurrently.
@@ 4 NeoSQuAD
The LLMs were chosen for diversity— they include some of the largest and some of the smallest models. The results are shown in Figure 1.333Note that we need to handle unparseable responses; this detail is discussed in Appendix A.
Figure 1: NeoSQuAD score curves across nine LLMs, normalized by the number of questions answered instead of the number of answers parsed. The shaded region around each curve represents ±$\pm 1\sigma$ confidence intervals.
Figure 1 allows us to observe several features of these models. Initially, we note that at $p=0$ (i.e., in the original evaluation), model performance is hard to distinguish. The accuracy is nearing saturation, and many models fall within the confidence intervals of others, making reliable differentiation difficult. However, at $p=0.3$, model performance has largely separated, reducing saturation and providing more informative scores due to the broader range of variation.
@@
The score at $p=0.3$ also provides a useful ordering of the larger models. In decreasing order of performance:
  • gemini-1.5-pro-002 and o1-preview (comparable)
  • gpt-4o-2024-08-06 and gemini-1.5-pro-001 (comparable)
  • Qwen2-72B-Instruct
@@
As discussed in Section 3, since there are three multiple-choice options, one might expect model performance to converge to $1/3$ at $p=1$. However, by $p=0.6$, the accuracy has dropped significantly below $1/3$ for all models except o1-preview.
[…]
Reflecting on the analysis, it appears that the most informative region of the score curve lies in $p\in[0.2,0.5]$. In this range, the garbling is sufficient to require reasoning, reduce saturation, and differentiate models, yet low enough to still observe meaningful performance differences.

### v3

@@ Abstract
As large language models (LLMs) become increasingly powerful, traditional evaluation metrics tend to saturate, making it challenging to distinguish between models. We propose a general method to transform existing LLM evaluations into a series of progressively more difficult tasks. These enhanced evaluations emphasize reasoning capabilities and can reveal relative performance differences that are not apparent in the original assessments.
To demonstrate the effectiveness of our approach, we create a new multiple-choice test corpus, extend it into a family of evaluations, and assess a collection of LLMs. Our results offer insights into the comparative abilities of these models, particularly highlighting the differences between base LLMs and more recent “reasoning” models.
@@ 1 Introduction
In this paper, we present a straightforward method for transforming any text-based evaluation into a family of increasingly challenging tests. This is a meta-evaluation technique: it takes an existing evaluation framework and converts it into a new one. To demonstrate our method, we generate a new evaluation framework and then extend it into a harder family of tests. Even this simple example reveals non-obvious differences in performance among models.
@@ 2 The Garbling Trick
Evaluations of language models have shifted from testing syntax (e.g., model perplexity) to assessing semantics (e.g., answering multiple-choice questions) [9]. The garbling trick enables us to interpolate between these extremes and examine both aspects simultaneously. As the garbling rate increases, the LLM faces two additional tasks not present in the original evaluation. First, the LLM must infer as much as possible about the underlying text from its garbled version. Second, given that it can only partially or uncertainly recover the text, the LLM must answer the question with incomplete information. Selecting the correct answer thus requires the LLM to engage in multiple forms of reasoning [^1] The term “reasoning” is used in this paper in two senses: to refer to the process by which any LLM selects an answer (as here), and to the subclass of LLMs that use test-time compute to improve inference. The meaning should be clear from context. concurrently.
@@ 4 NeoSQuAD Evaluation
5 NeoSQuAD Results
We applied the NeoSQuAD evaluation to 12 LLMs. The models were chosen for diversity— they include some of the largest and some of the smallest models. The models can be divided into 8 base LLMs and 4 “reasoning” LLMs (which leverage test-time compute). The traditional LLMs are Qwen2-72B-Instruct, Meta-Llama-3.1-70B-Instruct, Meta-Llama-3.1-8B-Instruct, Phi-3-medium-4k-instruct, Mixtral-8x7B-Instruct-v0.1, gemini-1.5-pro-001, gemini-1.5-pro-002, gpt-4o-2024-08-06. The reasoning LLMs are gemini-2.5-pro-preview-03-25, o1-preview-2024-09-12, o3-mini-2025-01-31, and Deepseek-R1. Note that the default reasoning effort was used for the reasoning models in all cases (e.g., o3-mini-2025-01-31 corresponds to o3-mini (medium)).
The results for traditional LLMs are shown in Figure 1.444Note that evaluator needs to handle unparseable responses; this detail is discussed in Appendix A. and the results for reasoning models are shown in Figure 2.
Figure 1: NeoSQuAD score curves across eight traditional (non-reasoning) LLMs, normalized by the number of questions answered instead of the number of answers parsed. The shaded region around each curve represents ±$\pm 1\sigma$ confidence intervals.
Figure 2: NeoSQuAD score curves across four reasoning LLMs, normalized by the number of questions answered instead of the number of answers parsed. The shaded region around each curve represents ±$\pm 1\sigma$ confidence intervals. The performance of the eight non-reasoning models is faintly shown in grey for comparison.
Figures 1 and 2 allow us to observe several general features of these models. At $p=0$ (i.e., in the original, ungarbled evaluation), model performance is hard to distinguish. The accuracy is nearing saturation, and many models fall within the confidence intervals of others, making reliable differentiation difficult. However, by $p=0.3$, model performance has largely separated, reducing saturation and providing more variable and hence more informative scores.
@@ 5 NeoSQuAD Results
The score at $p=0.4$ provides a useful ordering of the models. In decreasing order of performance:
  • gemini-2.5-pro-preview-03-25
  • o1-preview
  • gemini-1.5-pro-001
  • Tied: gpt-4o-2024-08-06, gemini-1.5-pro-001
  • Tied: Deepseek-R1, Qwen2-72B-Instruct
@@
  • o3-mini-2025-01-31
  • Tied: Phi-3-medium-4k-instruct, Mixtral-8x7B-Instruct-v0.1, Llama-3.1-8B-Instruct
@@
As discussed in Section 3, since there are three multiple-choice options, one might expect model performance to converge to $1/3$ at $p=1$. However, by $p=0.6$, the accuracy has dropped significantly below $1/3$ for all models except o1-preview and gemini-2.5-pro-preview-03-25.
If we were to draw some rules of thumb from this analysis, it appears that the most informative region of the score curve lies in $p\in[0.2,0.5]$, and particularly $p\in[0.3,0.4]$. In this range, the garbling is sufficient to require reasoning, reduce saturation, and differentiate models, yet low enough to still observe meaningful performance differences.
Finally, these score curves provide insights into non-reasoning versus reasoning models. The two largest reasoning models, o1-preview and gemini-2.5-pro-preview-03-25, are significantly better than all other models from $p=0.3$ onward. Note that o1-preview uses a gpt-4o base model, so the improvement granted by test-time compute is particularly clear. Note also that these were the only models that achieved a score of approximately $1/3$ accuracy in the high garble-rate regime. In other words, they were able to overcome the implicit bias towards incorrect answers that all the other models suffered from.
These results also reveal two distinct clusters among the reasoning models under increasing garbling. In particular, OpenAI’s o1-preview and o3-mini exhibit markedly different behavior: o1-preview remains robust even at high corruption rates, while o3-mini degrades nearly in step with non-reasoning baselines. Likewise, Google’s Gemini-2.5-Pro joins o1-preview in the top tier, whereas DeepSeek’s R1 tracks with the weaker group.
One possible explanation is model capacity, i.e., the number of active parameters engaged during inference. Because OpenAI and Google treat model parameters as proprietary, the true number of active parameters is not publicly known, so the following suggestions should be treated as highly speculative. Public reports suggest that o1-preview employs around 300 billion active parameters from a presumably much larger model [2]. In contrast, o3-mini is a 200 billion-parameter dense model [19]. Despite a total of 671 billion parameters in its Mixture-of-Experts architecture, DeepSeek’s R1 activates only about 37 billion at a time. Earlier Gemini releases reportedly used 1.56 trillion total parameters, so Gemini-2.5-Pro likely maintains a similarly large active footprint [20]. Perhaps, then, the poorer performance of o3-mini and R1 may be related to their seemingly more limited active parameter count.